# Evaluation of cosmogenic Ge-68 background in a high purity germanium detector via a time series fitting method

### v1

@@ Abstract
Ge-68 is a cosmogenic isotope in germanium with a half-life of 270.9 days. Ge-68 and its decay daughter Ga-68 contribute considerable background with energy up to 3 MeV to low background $\gamma$ spectrometers using high purity germanium (HPGe) detectors. In this paper, we evaluated the background of Ge-68 and Ga-68 in a $p$-type coaxial HPGe detector operated at China Jinping underground laboratory (CJPL) via a time series fitting method. Under the assumption that Ge-68 and Ga-68 are in radioactive equilibrium and airborne radon daughters are uniformly distributed in the measurement chamber of the spectrometer, we fit the time series of count rate in 1-3 MeV to calculate the Ge-68 activity, radon daughter concentrations, and the time-invariant background component. Total 90 days measured data were used in analysis, a hypothesis test confirmed a significant Ge-68 signal at 99.64% confidence level. The initial activity of Ge-68 is fitted to be 477.0$\pm$112.4 $\mu$Bq/kg, corresponding to an integral count rate of 55.9 count/day in 1-3 MeV range. During the measurement, Ge-68 activity decreased by about 30%, contributing about 62% of the total background in 1-3 MeV range. Our method also provides an estimation of the variation of airborne radon daughter concentrations in the measurement chamber, which could be used to monitor the performance of radon reduction measures.
@@ 1 Introduction
High purity germanium (HPGe) detectors have been widely used in radiation monitoring, nuclear physics, particle physics, and astrophysics due to its high energy reslution, high stopping power and low intrinsic background [1, 2, 3]. The use of HPGe detector as $\gamma$ spectrometer requires strict control and accurate meausurement of background for detecting trace radioactivity within the sample. For $\gamma$ spectrometers operated at underground laboratorys, the rock overburden reduces the cosmic-ray muon flux by several orders of magnitudes, for instance, 1000 m rock can reduce cosmic-ray muon flux by over $10^{5}$ times [4]. External $\gamma$ can be shielded by high-Z materials like copper and lead, neutron can be shielded by borated polyethylene. However, the cosmogenic radioactive isotopes cumulated during the fabrication and transportation of detector above ground can continuously contribute background after the detector have been moved underground. As the production of cosmogenic isotopes becomes negligible underground [5], their activities decrease after detector arrived underground, leading to a change of background in the measured spectrum. Therefore, a measurement of their initial activity and an evaluation of their background during detector’s underground operation are important for the understanding and modeling of the background.
@@
Figure 1: Left panel: decay scheme of Ge-68 and Ga-68, data from[6]. Right panel: typical Ge-68 and Ga-68 spectra in a HPGe detector, the spectra are simulated via Geant4[7, 8, 9] using a 1-kg cylinder Ge model with 0.5 mm deadlayer at the surface (energy resolution not considered).
@@
Ge-68 is a cosmogenic isotope produced by nuclear reactions between germanium nucleus and high energy cosmic-rays (neutron, proton, and $\gamma$). At sea level, the production rate is around 80-120 $\rm kg^{-1}day^{-1}$ for different latitudes and longitudes [10], corresponding to a saturated activity of 0.93-1.39 mBq/kg. Ge-68 decays to Ga-68 via electron capture (EC) with a half-life of 270.9 days (decay scheme in Fig.1) [6]. Its decay emits X-rays and auger electrons with total energy equals to the binding energy of the captured electron ($<$11 keV), therefore only contributes background at low energy region (as in Fig.1(right)). Ga-68 decays to Zn-68 with a $Q$-value of 2921.1 keV and a half-life of 67.7 min [6], its decay scheme is shown in Fig.1(left). In most case, Ga-68 decays to ground state of Zn-68 via $\beta^{+}$ mode and emit a positron with maximum kinetic energy up to 1899.1 keV. The spectrum shape of Ga-68 in HPGe detector is shown in Fig.1(right), it manifests a arc-shaped line with a small peak from 511 keV annihilation photons and summed peaks around 1 MeV from $\gamma$ and X-rays in EC decay mode. Due to the relative short half-life of Ga-68, it will be in radioactive equilibrium with Ge-68. Ga-68 with activity of 1 mBq/kg could contribute a background of approximate 85 count/day/kgGe at 60-2700 keV range, which is comparable to the typical background level of an underground HPGe $\gamma$-spectrometer [11].
@@
There are general two ways of measuring the Ge-68 activity, one is analyzing the 10.38 keV KX-ray peak in the spectrum [10, 12, 13]. However, this method is not feasible for detectors with dynamical range cannot cover the low energy range. The other way is fitting the count rate at different time. Once the detector arrives underground, the decay of Ge-68 and Ga-68 will lead to a decrease of count rate in the spectrum. Fitting the count rate with Ge-68 half-life could provide an estimation of the activity, but requires decoupling other time variant components, e.g., count rate change caused by concentration variation of the airborne radon daughter. We develop a method fitting the count rate with Ge-68 decay component and the variation of radon daughter simultaneously while constraining the radon daughter concentration by its characteristic peaks. This method is then applied to measuring the activity of Ge-68 in a HPGe detector operated in China Jinping underground laboratory (CJPL)[14].
@@ 2.1 HPGe detector at CJPL
The HPGe detector studied is a low background $p$-type coaxial HPGe detector purchase from CANBERRA, and is used as the detector of a low background $\gamma$-spectrometer at CJPL. It has a Ge crystal of 2.48 kg with 0.5 mm deadlayer and a energy threshold of 100 keV, the crystal and its surrounding structure are shown in the left panel of Fig.2 [15]. After the detector been manufactured in Fracnce, it was shipped to CJPL via truck and train, it arrived at CJPL at 2020/07/25 with approximate 1 month exposre time above ground [15].
@@
After arrived at CJPL, the detector is shielded by copper and lead to reduce the background from enviromental $\gamma$-rays. The shielding is made of 10 cm underground storage copper and 20 cm lead (5 cm ancient Poland lead and 15 cm modern lead), the structure of the Cu/Pb shielding is illustrated at Fig.2. Nitrogen gas has been constantly injected into the detector chamber inside copper shielding to reduce the radon concentration.
The energy calibration was performed using Co-60, Co-57, and Eu-152 sources[15]. After the calibration, the sources are removed and detector was operated to measure the background. The spectrum was saved regularly, each saved spectrum corresponds to 6 hours live measurement time. Total 90 days background measurement data is cumulated and used to calculate the activities of cosmogenic isotopes. The measurement of Mn-54, Co-57, and Co-58 activities via fitting their characteristic peaks could be found in[15]., this work focus on the measurement of Ge-68 activity.
@@ 2.2 Time series fitting method
The calculation of Ge-68 activity via fitting the variation of count rate requires decoupling various time variant and invariant background. The time-variant background considered are cosmogenic isotopes with relative short half-life and radon daughters in the void volume between the detector and copper shield. The time invariant background are those from radioisotopes with long half-life, for instance, $\gamma$-rays from U, Th decay chain and K-40. And for background from Co-60, as the half-life of Co-60 (1925.3 day) is much longer than the measure time (90 days), its activity decreases only 3% during the measurement, therefore, it is treated as a time invariant background in this work. The count rate in a specified energy range can be modeled as eq.2.2
@@
the specified energy range is indicated by $i$, $R_{i}(t,t+\Delta t)$ is the count rate in $i$ energy range at $t\sim t+\Delta t$ interval. The first item in eq.2.2 is the sum of contribution from each cosmogenic isotope $k$, $A_{0,k}$ is the initial specific activity at $t=0$ (units: Bq/kg), and $T_{1/2,k}$ is the decay half-life. $\varepsilon_{k,i}$ is the detection efficiency of the decay produts of $k$ isotope in $i$ energy range (units: cpd/(Bq/kg), cpd as count per day). $N_{Iso}$ is the total number of cosmogenic isotopes. The second item $\sum_{p}C_{p,i}\cdot\varepsilon_{p,i}$ is the sum of contribution from radon daughters. In this work, we treat each radon daughter independently as they are not in equilibrium, but make the assumption that they are uniformly distributed in the void volume inside copper shield. $C_{p}(t)$ is the average concentration of radon daughter $p$ (units: Bq/m3) in $t\sim t+\Delta t$ interval. $\varepsilon_{p,i}$ is the detection efficiency of $\gamma$-rays emitted by $p$ isotope in $i$ energy range (units: cpd/(Bq/m3)). $N_{Rn}$ is the total number of considered radon daughter. $B_{i}$ is the time-invariant component in $i$ energy range (units: cpd).
We built the detector and its shielding structure in Geant4[7, 8, 9] to simulate the detection efficiency of different background sources. The simulation is performed using Geant4 version 11.0.3 with the $shielding$ physics list. In simulation, the energy deposition in the surface deadlayer is not recorded, and the energy resolution of the detector is considered using the calibrated energy reslution function: $FWHM/\mathrm{keV}=0.9903+0.3197\sqrt{E/\mathrm{keV}+5.789\times 10^{-5}\times(E/\mathrm{keV})^{2}}$ [16]. The efficiencies $\varepsilon_{k,i}$ and $\varepsilon_{p,i}$ are then calculated by:
@@
where $N_{S}$ is the total number of simulated particles, $n_{i}$ is the observed count in $i$ energy range in simulated spectrum, $m_{Ge}$=2.48 kg is the mass of Ge crystal, $V_{Chamber}$=9.4 dm3 is the volume of the detector chamber inside copper shielding.
The activity of cosmogenic isotope ($A_{0,k}$) and radon daughter concentration ($C_{p}(t)$) can be calculated via fitting the modeled count rate to measured data using a maximum likelihood method. The observed count in each time interval obeys Possion distribution, therefore the likelihood function can be written as:
@@
where $\lambda_{i,t}$ and $n_{i,t}$ is the expected and observed count in $i$ energy range at $t\sim t+\Delta t$ interval. $N_{E}$ is the numbere of selected energy interval and $N_{T}$ is the numbere of time interval. The total number of fitting parameters is $N_{Iso}+N_{Rn}\cdot N_{T}+N_{E}$, corresponding to number of cosmogenic isotopes, radon daughter concentration, and time-invariant background. And the number of observed data is $N_{E}\cdot N_{T}$ corresponding to the count rate seriess in each energy range. To better constraine the fitting parameters, multiple energy ranges are required, they should at least include the signal of cosmogenic isotopes and characteristic peak of radon daughter. It should be noted that the fitted of radon daughter concentration $C_{p}(t)$ is only sensitive to the variation of its concentration, the constant part will be regarded as a time-invariant background.
In the calculation of Ge-68 activity, we select three energy ranges: 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV, the first two are characteristic peaks of radon daughter Bi-214, and the third is choosed as the signal region of Ge-68 and its decay daughter Ga-68. We assume the Ge-68 and Ga-68 are in equilibrium, and hereafter use Ge-68 to indicate Ge-68 and Ga-68. We only considere two time variant components: Ge-68 and radon daughter Bi-214. Using activities determined in our previous work[15], other cosmogenic isotopes (Mn-54, Co-57, and Co-58) contributes negligible background in the selected energy regions (less than 1%), therefore are omitted in this work. The total 358 spectra measured in 6 hours interval are merged into 30 spectra, each corresponds to 3 days live measurement time. The count rate in each energy range is calculate for the 30 spectra, and total 90 count rate data are used for analysis.
The fitting of the parameters is by maximum the likelihood function $\mathcal{L}$ in eq.4. We use the Markov Chain Monte Carlo (MCMC) method[20] to calculate the best fit result and the corresponding uncertainty. The calculation is performed using the $UltraNest$ MCMC toolkit[17] in $python3$ platform.
@@ 2.3 Statistic test of the significance of Ge-68 signal
In order to evaluate the significance of Ge-68 signal, we perform a hypothesis test: the null hypothesis ($H_{0}$) is that there are no Ge-68 in the measured data, and the alternative hypothesis ($H_{1}$) is the best fit result. For the null hypothesis, Bi-214 concentration and time-invariant background are fitted using the same procedure while setting the Ge-68 activity to 0.
The $P$ value is used to test the coincidence between the hypothesis and measurement. It is defined as the probability of getting a worse result than observed in the measurement under a specified hypothesis[18]. In this work, it is calculated by:
@@
where $s(H_{i})$ is the test statistic defined as the sum of likelihood value in the 1-3 MeV Ge-68 signal region. $\lambda_{t}(H_{i})$ is the expected count in $t\sim t+\Delta t$ interval under $H_{i}$ hypothesis. $s_{obs}(H_{i})$ is the observed value of $s(H_{i})$, and is calculated via eq.7 using measured data. $f(s|H_{i})$ is the probability distribution function (PDF) of test statistic $s$ under $H_{i}$ hypothesis.
$f(s|H_{i})$ is calculated via a toy Monte Carlo method: $n_{t}$ are randomly sample by Possion distribution using $\lambda_{t}(H_{i})$ as the expectation. $s(H_{i})$ is then calculated and storaged for a group of sampled $n_{t}$ ($t=1,2,...,N_{T}$) as a hypothesis experiment. Then the hypothesis experiment is performed for 20,000 times to get the PDF of $s(H_{i})$.
The $P$ value of the alternative hypothesis ($P(H_{1})$) demonstrates the goodness of the fit, while the $P$ value of the null hypothesis ($P(H_{0})$) indicates if adding Ge-68 in fitting is necessary. The significance of Ge-68 signal is the combination of a small $P(H_{0})$ and a large $P(H_{1})$.

### v2

@@ Abstract
Ge-68 is a cosmogenic isotope in germanium with a half-life of 270.9 days. Ge-68 and its decay daughter Ga-68 contribute considerable background with energy up to 3 MeV to low background $\gamma$ spectrometers using high purity germanium (HPGe) detectors. In this paper, we evaluated the background of Ge-68 and Ga-68 in a $p$-type coaxial HPGe detector operated at China Jinping underground laboratory (CJPL) via a time series fitting method. Under the assumption that Ge-68 and Ga-68 are in radioactive equilibrium and airborne radon daughters are uniformly distributed in the measurement chamber of the spectrometer, we fit the time series of count rate in 1-3 MeV to calculate the Ge-68 activity, radon daughter concentrations, and the time-invariant background component. A total of 90-day measurement data were used in the analysis, a hypothesis test confirmed a significant Ge-68 signal at 99.64% confidence level. The initial activity of Ge-68 is fitted to be 477.0$\pm$112.4 $\mu$Bq/kg, corresponding to an integral count rate of 55.9 count/day in the 1-3 MeV range. During the measurement, Ge-68 activity decreased by about 30%, contributing about 62% of the total background in the 1-3 MeV range. Our method also provides an estimation of the variation of airborne radon daughter concentrations in the measurement chamber, which could be used to monitor the performance of radon reduction measures.
@@ 1 Introduction
High purity germanium (HPGe) detectors have been widely used in radiation monitoring, nuclear physics, particle physics, and astrophysics due to its high energy resolution, high stopping power, and low intrinsic background [1, 2, 3]. The use of an HPGe detector as $\gamma$ spectrometer requires strict control and accurate measurement of background for detecting trace radioactivity within the sample. For $\gamma$ spectrometers operated at underground laboratories, the rock overburden reduces the cosmic-ray muon flux by several orders of magnitudes, for instance, 1000 m rock can reduce cosmic-ray muon flux by over $10^{5}$ times [4]. External $\gamma$ can be shielded by high-Z materials like copper and lead, neutron can be shielded by borated polyethylene. However, the cosmogenic radioactive isotopes cumulated during the fabrication and transportation of the detector above ground can continuously contribute to the background after the detector has been moved underground. As the production of cosmogenic isotopes becomes negligible underground [5], their activities decrease after the detector arrives underground, leading to a change of background in the measured spectrum. Therefore, measuring the cosmogenic isotope activities and evaluating their background in an underground HPGe spectrometer is important for the background modeling and understanding.
@@
Figure 1: Left panel: decay scheme of Ge-68 and Ga-68, data from[6]. Right panel: typical Ge-68 and Ga-68 spectra in an HPGe detector, the spectra are simulated via Geant4[7, 8, 9] using a 1-kg cylinder Ge model with 0.5 mm dead layer at the surface (energy resolution not considered).
@@
Ge-68 is a cosmogenic isotope produced by nuclear reactions between the germanium nucleus and high energy cosmic rays (neutron, proton, and $\gamma$). At sea level, the production rate is around 80-120 $\rm kg^{-1}day^{-1}$ for different latitudes and longitudes [10], corresponding to a saturated activity of 0.93-1.39 mBq/kg. Ge-68 decays to Ga-68 via electron capture (EC) with a half-life of 270.9 days (decay scheme in Fig.1) [6]. Its decay emits X-rays and auger electrons with total energy equaling to the binding energy of the captured electron ($<$11 keV), therefore only contributing background at low energy region (as in Fig.1(right)). Ga-68 decays to Zn-68 with a $Q$-value of 2921.1 keV and a half-life of 67.7 min [6], its decay scheme is shown in Fig.1(left). In most cases, Ga-68 decays to the ground state of Zn-68 via $\beta^{+}$ mode and emits a positron with a maximum kinetic energy up to 1899.1 keV. The spectrum shape of Ga-68 in a cylindrical HPGe detector is shown in Fig.1(right), it manifests an arc-shaped line with a small peak from 511 keV annihilation photons and summed peaks around 1 MeV from $\gamma$ and X-rays in the EC decay mode. Due to the relatively short half-life of Ga-68, it will be in radioactive equilibrium with Ge-68. Ga-68 with an activity of 1 mBq/kg could contribute a background of approximately 85 count/day/kgGe at 60-2700 keV range, which is comparable to the typical background level of an underground HPGe $\gamma$-spectrometer [11]. As the energy of Ge-68 background spans from keV to MeV, it is also an important background both in dark matter direct search experiments (SuperCDMS[12] and CDEX[13]) and neutrinoless double beta decay experiments (GERDA[14], Majorana[15], LEGEND[16]). The measurement of the Ge-68 activity is important for the background understanding of those low background experiments[10, 17, 18].
@@
There are general two ways of measuring the Ge-68 activity, one is analyzing the 10.38 keV KX-ray peak in the spectrum [10, 19, 20]. However, this method is not feasible for detectors with a dynamic range that cannot cover the low energy range. The other way is fitting the count rate at different times. Once the detector arrives underground, the decay of Ge-68 and Ga-68 will decrease the count rate in the spectrum. Fitting the count rate with Ge-68 half-life could provide an estimation of the activity, but requires decoupling other time variant components, e.g., the count rate change caused by concentration variation of the airborne radon daughter. We develop a method fitting the count rate with the Ge-68 decay component and the variation of radon daughter simultaneously while constraining the radon daughter concentration by its characteristic peaks. This method is then applied to measuring the activity of Ge-68 in an HPGe detector operated in China Jinping underground laboratory (CJPL)[21].
@@ 2.1 HPGe detector at CJPL
The HPGe detector studied is a low background $p$-type coaxial HPGe detector purchased from CANBERRA and is used as the detector of a low background $\gamma$-spectrometer at CJPL. It has a Ge crystal of 2.48 kg with 0.5 mm dead layer and an energy threshold of 100 keV, the crystal and its surrounding structure are shown in the left panel of Fig.2 [22]. After the detector had been manufactured in France, it was shipped to CJPL via truck and train, it arrived at CJPL on 2020/07/25 with an approximate 1 month exposure time above ground [22].
@@
After arriving at CJPL, the detector is shielded by copper and lead to reduce the background from environmental $\gamma$-rays. The shielding is made of 10 cm underground storage copper and 20 cm lead (5 cm ancient Poland lead and 15 cm modern lead), the structure of the Cu/Pb shielding is illustrated in Fig.2. Nitrogen gas has been constantly injected into the detector chamber inside the copper shielding to reduce the radon concentration.
The energy calibration was performed using Co-60, Co-57, and Eu-152[22]. After the calibration, the sources were removed and the detector was operated to measure the background. The spectrum was saved regularly, each saved spectrum corresponds to 6 hours of live measurement time. The data taken started in 2020/08/15 and ended in 2020/11/21, a total of 90 days live time background measurement data is cumulated and used to calculate the activities of cosmogenic isotopes. The measurement of Mn-54, Co-57, and Co-58 activities via fitting their characteristic peaks could be found in[22]. This work focuses on the measurement of Ge-68 activity.
@@ 2.2 Time series fitting method
The calculation of Ge-68 activity via fitting the variation of count rate requires decoupling various time variant and invariant background. The time-variant backgrounds considered are cosmogenic isotopes with relatively short half-lives and radon daughters in the void volume between the detector and copper shield. The time-invariant background are those from radioisotopes with long half-lives, for instance, $\gamma$-rays from U, Th decay chain, and K-40. And for background from Co-60, as the half-life of Co-60 (1925.3 days) is much longer than the measurement time (90 days), its activity decreases only 3% during the measurement. Therefore, it is treated as a time-invariant background in this work. The count rate in a specified energy range can be modeled as Eq.2.2
@@
the specified energy range is indicated by $i$, $R_{i}(t,t+\Delta t)$ is the count rate in $i$ energy range at $t\sim t+\Delta t$ interval. The first item in Eq.2.2 is the sum of the contribution from each cosmogenic isotope $k$, $A_{0,k}$ is the initial specific activity at $t=0$ (units: Bq/kg), and $T_{1/2,k}$ is the decay half-life. $\varepsilon_{k,i}$ is the detection efficiency of the decay produts of $k$ isotope in $i$ energy range (units: cpd/(Bq/kg), cpd as count per day). $N_{Iso}$ is the total number of cosmogenic isotopes. The second item $\sum_{p}C_{p,i}\cdot\varepsilon_{p,i}$ is the sum of the contribution from radon daughters. In this work, we treat each radon daughter independently as they are not in equilibrium, but make the assumption that they are uniformly distributed in the measurement chamber. $C_{p}(t)$ is the average concentration of radon daughter $p$ (units: Bq/m3) in $t\sim t+\Delta t$ interval. $\varepsilon_{p,i}$ is the detection efficiency of $\gamma$-rays emitted by $p$ isotope in $i$ energy range (units: cpd/(Bq/m3)). $N_{Rn}$ is the total number of considered radon daughters. $B_{i}$ is the time-invariant component in $i$ energy range (units: cpd).
We built the detector and its shielding structure in Geant4[7, 8, 9] to simulate the detection efficiency of different background sources. The geometry in Geant4 includes the crystal, its holding structure, the cryostat, and the shielding. The simulation is performed using Geant4 version 11.0.3 with the $shielding$ physics list. In the simulation, the energy deposition in the surface dead layer is not recorded, and the energy resolution of the detector is considered using the calibrated resolution function: $FWHM/\mathrm{keV}=0.9903+0.3197\sqrt{E/\mathrm{keV}+5.789\times 10^{-5}\times(E/\mathrm{keV})^{2}}$ [23]. The efficiencies $\varepsilon_{k,i}$ and $\varepsilon_{p,i}$ are then calculated by:
@@
where $N_{S}$ is the total number of simulated particles, $n_{i}$ is the observed count in $i$ energy range in the simulated spectrum, $m_{Ge}$=2.48 kg is the mass of Ge crystal, $V_{Chamber}$=9.4 dm3 is the volume of the detector chamber inside the copper shielding.
The activity of cosmogenic isotope ($A_{0,k}$) and radon daughter concentration ($C_{p}(t)$) can be calculated by fitting the modeled count rate to measured data using a maximum likelihood method. The observed count in each time interval obeys the Possion distribution, and the likelihood function can be written as:
@@
where $\lambda_{i,t}$ and $n_{i,t}$ is the expected and observed count in $i$ energy range at $t\sim t+\Delta t$ interval. $N_{E}$ is the number of selected energy intervals, and $N_{T}$ is the number of time intervals. The total number of fitting parameters is $N_{Iso}+N_{Rn}\cdot N_{T}+N_{E}$, corresponding to the number of cosmogenic isotopes, radon daughter concentration, and time-invariant background. And $N_{E}\cdot N_{T}$ is the number of observed data, corresponding to the count rate series in each energy range. To better constrain the fitting parameters, multiple energy ranges are required, they should at least include the signal of cosmogenic isotopes and characteristic peak of radon daughter. It should be noted that the fitting of the radon daughter concentration $C_{p}(t)$ is only sensitive to the variation of its concentration, the constant part will be regarded as a time-invariant background.
In the calculation of Ge-68 activity, we select three energy ranges: 609$\pm$5 keV, 1764$\pm$6 keV, and 1000$\sim$3000 keV, the first two are characteristic peaks of the radon daughter Bi-214, and the third is chosen as the signal region of Ge-68 and its decay daughter Ga-68. We assume the Ge-68 and Ga-68 are in equilibrium, and hereafter use Ge-68 to indicate Ge-68 and Ga-68. We only consider two time variant-components: Ge-68 and the airborne radon daughter Bi-214. Based on activities determined in our previous work[22], other cosmogenic isotopes (Mn-54, Co-57, and Co-58) contribute negligible background in the selected energy regions (less than 1%), therefore are omitted in this work. The total 358 spectra measured in a 6-hour interval are merged into 30 spectra, each corresponding to 3 days of live measurement time. The count rate in each energy range is calculated for the 30 spectra, and a total of 90 count rate data are used for analysis.
The fitting of the parameters is by maximum the likelihood function $\mathcal{L}$ in Eq.4. We use the Markov Chain Monte Carlo (MCMC) method[24] to calculate the best-fit result and the corresponding uncertainty. The calculation is performed using the $UltraNest$ MCMC toolkit[25] in the $python3$ platform.
@@ 2.3 Statistic test of the significance of Ge-68 signal
In order to evaluate the significance of the Ge-68 signal, we perform a hypothesis test: the null hypothesis ($H_{0}$) is that there are no Ge-68 in the measured data, and the alternative hypothesis ($H_{1}$) is the best-fit result. For the null hypothesis, Bi-214 concentration and time-invariant background are fitted using the same procedure while setting the Ge-68 activity to 0.
The $P$ value is used to test the consistency between the hypothesis and measurement. It is defined as the probability of getting a worse result than observed in the measurement under a specified hypothesis[26]. In this work, it is calculated by:
@@
where $s(H_{i})$ is the test statistic defined as the sum of the logarithmic likelihood value in the 1-3 MeV Ge-68 signal region. $\lambda_{t}(H_{i})$ is the expected count in $t\sim t+\Delta t$ interval under $H_{i}$ hypothesis. $s_{obs}(H_{i})$ is the observed value of $s(H_{i})$, and is calculated via Eq.7 using measured data. $f(s|H_{i})$ is the probability distribution function (PDF) of test statistic $s$ under the $H_{i}$ hypothesis.
$f(s|H_{i})$ is calculated via a toy Monte Carlo method: $n_{t}$ are randomly sample by Possion distribution using $\lambda_{t}(H_{i})$ as the expectation. $s(H_{i})$ is then calculated and stored for a group of sampled $n_{t}$ ($t=1,2,...,N_{T}$) as a hypothesis experiment. Then the hypothesis experiment is performed 20,000 times to get the PDF of $s(H_{i})$.
The $P$ value of the alternative hypothesis ($P(H_{1})$) demonstrates the goodness of the fit, while the $P$ value of the null hypothesis ($P(H_{0})$) indicates if adding Ge-68 in fitting is necessary. The significance of the Ge-68 signal is the combination of a small $P(H_{0})$ and a large $P(H_{1})$.